Rollie is 7 years older than Pollie. 3 years ago, rollie was twice as old as pollie. find their present ages.
Answer by htmentor(1343)

Let R = Rollie's present age
Then Pollie's age is R - 7
Given: 3 years ago, Rollie was two times as old as Pollie, so we can write
R-3 = 2((R-7)-3)
R-3 = 2R-20
This gives R=17
So Rollie's age = 17 and Pollies age = 17 - 7 = 10
Check:
3 years ago their ages were 14 and 7:
14 = 2*7